OpenStudy (anonymous):

A 10.00 g sample of a compound containing C, H, and O is burned completely to produce 14.67 g of CO2 and 6.000 g of H2O. What is the empirical formula of this compound? (A) CHO (B) CH2O (C) CH2O2 (D) C2H4O

OpenStudy (cuanchi):

All the C in the CO2 is coming from the C in the compound All the H in the water is coming from the H in the compound The amount of Oxygen in the compound you calculate by difference. You can do the calculation in grams or in moles 1. calculate how many grams of C are in 14.67g of CO2 14.67 g CO2 x 12 g C / 44 g CO2 = ??? g C 14.67 g CO2 x 1mol C / 44 g CO2 = ??? mol C 2. calculate how many grams of H are in 6.000g of H2O 6.000g of H2O x 18 g H2O / 44 g CO2 = ??? g H 6.000g of H2O x 1mol H2O/ 44 g CO2 = ??? mol H 3. gC + gH + gO = 10g sample compound gO = 10g compound - gC -gH 4. convert the g O to mole of O g O x 1mole O / 16g O = ???? mole O 5. build a pseudoformula with the moles of C, O, H Cx Oy Hz divide x, y and z by the smallest of them 6. adjust the values to whole numbers depending on the decimal part multiplying by Common Decimal Equivalent Fraction Mole Ratio Example 0.125 1/8 1 : 1.125 converts to 1 : 9/8 multiply throughout by 8 to give 8 : 9 0.25 1/4 1 : 0.25 converts to 1 : 1/4 multiply throughout by 4 to give 4 : 1 0.33 1/3 1 : 1.33 converts to 1 : 4/3 multiple throughout by 3 to give 3 : 4 0.375 3/8 1 : 1.375 converts to 1 : 11/8 multiply throughout by 8 to give 8 : 11 0.5 1/2 2 : 1.5 converts to 2 : 3/2 multiply throughout by 2 to give 4 : 3 0.625 5/8 1 : 1.625 converts to 1 : 13/8 multiply throughout by 8 to give 8 : 13 0.667 2/3 2 : 1.66 converts to 2 : 5/3 multiply throughout by 3 to give 6 : 5 0.875 7/8 1 : 0.875 converts to 1 : 7/8 multiply throughout by 8 to give 8 : 7

OpenStudy (cuanchi):

http://www.ausetute.com.au/empirical.html

OpenStudy (anonymous):

thank you dear

OpenStudy (anonymous):

6.000g of H2O x 18 g H2O / 44 g CO2 should I divided by 44 or 18 b\c it is for the water

OpenStudy (cuanchi):

yes by 2/18 sorry I copy the same from above and I forgot to correct the mass of the CO2 by the mass of H2O 2. calculate how many grams of H are in 6.000g of H2O 6.000g of H2O x 2 g H / 18 g H2O= ??? g H 6.000g of H2O x 2mol H/ 18 g H2O = ??? mol H